OpenStudy (bookworm14):

@satellite73 @hilbertboy96 i could use some help , if you have enough time of course.... (i will post it below)

OpenStudy (anonymous):

I got time

OpenStudy (anonymous):

Well what you want to do is isolate the variable as much as you can that you need to solve so on the left since j is on the left, 7k, 10l and 4 have nothing to do with j so move them to the other side and you got 8jkl= whatever you get from the first step. You then divide by everything else but j to get j on its own.

OpenStudy (anonymous):

don't you hate math?

OpenStudy (bookworm14):

so what would the answer be (sry i usually try to get some visuals in my mind) @satellite73 I do not hate math, i am usually extremely good at it, but i just don't "enjoy" math its too stressful for my liking

OpenStudy (anonymous):

\[\huge 9\color{red}jkl +7k+10l+4=-9k-3l+8\]

OpenStudy (anonymous):

Well as i said when you move that stuff to the other side you get.... \[8jkl = -16k - 13l + 4\] Divide by 8kl and you get \[j = \frac{ -16k -13l +4 }{ 8kl }\] (First choice)

OpenStudy (anonymous):

subtract \(7k\) and subtract \(10l\) and subtract \(4\) from the left hand side

OpenStudy (anonymous):

and you get what @hilbertboy96 wrote

OpenStudy (anonymous):

Yup. I am just a guy who can solve math in his head quick

OpenStudy (bookworm14):

I kind of get it... (sorry im just so new to this, ive never struggled like this before hah)

OpenStudy (bookworm14):

here is another one (i have a few) i will try to solve it also and then we can compare answers k? if i get stuck i will tell you

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Just let me know if you get stuck

OpenStudy (bookworm14):

ughhhh im lost x_X

OpenStudy (anonymous):

Haha ok. I wrote this one out for myself so I will show and explain the steps. We have..... \[6fg + 6fh - 4f -8 = -9g + 8\] This is a tougher one to solve for f compared to the last one since there are more then one term with f in it correct?

OpenStudy (bookworm14):

yes, i got lost lol

OpenStudy (anonymous):

That is fine lol. OK we got one term that has no variable so that moves to the right and we get.... \[6fg + 6fh - 4f = -9g + 16\] Can you think of what we might need to do next?

OpenStudy (anonymous):

We got three terms and each one has f in common so what can we do knowing this?

OpenStudy (anonymous):

Or is this what is confusing and losing you?

OpenStudy (bookworm14):

im just lost all together haha uhmm would we carry the 6g and 6h to the other side ??????

OpenStudy (anonymous):

we can not do that since the f would go with it and then we got f on both sides. We can factor f out like so.... \[f(6g + 6h -4) = -9g + 16\] You see how I did that and why I can?

OpenStudy (bookworm14):

yes i understand how you did that

OpenStudy (anonymous):

ok now we divided everything that is multiplied by f over and we solve f.... \[f = \frac{ -9g + 16 }{ 6g + 6h - 4 }\] This is one of your choices correct?

OpenStudy (bookworm14):

yess, wow that actual seemed really simple lol xD , are you busy i have a few more?

OpenStudy (anonymous):

I am fairly unbusy and bored. So throw what you got at me

OpenStudy (bookworm14):

i got textbooks, and pencils, and notepads should i go on and throw those at you ? lol

OpenStudy (anonymous):

Sure! lol

OpenStudy (bookworm14):

Yess!! lol here is the prob i will try to solve also

OpenStudy (anonymous):

Shall give you hint. Follow same format as last problem

OpenStudy (bookworm14):

okay i will

OpenStudy (bookworm14):

p(-6q+5r-10)= 6q-9 ?? am i on the right track or no?

OpenStudy (anonymous):

nope you need to do one step first then factor. Add the term without the p to the other side to get -6pq + 5pr -p = 6q + 1 Then factor p out of the left and divide

OpenStudy (bookworm14):

so i should do your step first : -6pq + 5pr -p = 6q + 10 then my step: p(-6q+5r-p)= 6q- 19 (or should i put "9+10 ??? ) equaling in ???

OpenStudy (anonymous):

Ok I will do as i did before by showing steps and explaining \[-6pq + 5pr - p +10 = 6q - 9\] We subtract ten to both sides to get \[-6pq + 5pr -p = 6q + 1 \] (since -9 + 10 = 1 on the right) Now we factor p out \[p(-6q + 5r -1) = 6q + 1\] (and notice the -1 on the left is because if you mutply the p back in that is how you get the -p) Now divide and get \[p = \frac{ 6q + 1 }{ -6q + 5r + 1 }\]

OpenStudy (anonymous):

Ahhh crud xD first step in subtracting 10 and also the factoring i messed up so final answer should be \[p = \frac{ 6q - 19 }{ -6q + 5r - 1 }\]

OpenStudy (anonymous):

Well nott the factoring just lost my minus when transfering my answer

OpenStudy (bookworm14):

okay yea i was gonna say something, but then again i ddn't want to confuse you since you know this way better than me lol

OpenStudy (anonymous):

Even the best people at math make simple mistakes

OpenStudy (bookworm14):

and you actually messed up again lol 9-10 = -1 so it would be -6q-1 (for the numerator) (and the site im on can prove it lol)

OpenStudy (anonymous):

no it is -19 since you got -9 and you are subtracting 10 so you have -9-10 = -19

OpenStudy (bookworm14):

your wya makes more sense but can i show you the explanation they gave me (it only shows me after I answer lol)

OpenStudy (anonymous):

ahhh ok i see what they are doing....ok let me redo my work for you lol

OpenStudy (anonymous):

and yes also i transferred the problem wrong xD

OpenStudy (bookworm14):

here is the mess they show me afterwards lol

OpenStudy (anonymous):

\[-6pq + 5pr - p - 10 = 6q - 9\] \[-6pq + 5pr - p = 6q + 1\] \[-p(6pq-5pr+1) = 6q + 1\] (the reason for -p is because they are preventing a negative in the denominator which is a rule that you can not have) \[-p = \frac{ 6q + 1 }{ 6q - 5r +1 }\] \[p = \frac{ -1(6q +1) }{ 6q - 5r + 1 }\] \[p = \frac{ -6q -1 }{ 6q -5r +1 }\]

OpenStudy (anonymous):

yea basically what I got lol

OpenStudy (anonymous):

but i just put the negative in the top a different way that works

OpenStudy (anonymous):

and in third step suppose to be no p's in the parenthesies xD

OpenStudy (bookworm14):

well that one was weird one more ques (of this topic) thank you so much for helping me!!!

OpenStudy (anonymous):

yep anytime!

OpenStudy (anonymous):

Ok this one I will help you through since the last ones were confusing

OpenStudy (bookworm14):

ohhh this is one of those weirdy ones ahhh, i wish to be an astronomer or cosmologist but i will never make it if i can't understand this math :'(

OpenStudy (anonymous):

You will make it. Practice is all you need and guidance. Ok first I will write out the original \[\frac{ -2r + 7s }{ 3 } = \frac{ 4r + 4t }{ 6 }\] We got two fractions here which is very unconventional! So first thing we want to do is get rid of these fractions by multiplying both sides by 6 since that will cancel the 6 out on the right and the 3 on the left as you will see.... \[\frac{ 6(-2r + 7s) }{ 3 } = \frac{ 6(4r + 4t) }{ 6 }\] The 6 over 6 on the right cancels but the 6 and 3 on the left makes the 3 disappear and the 6 becomes a 2 since you are dividing it by 3..... \[2(-2r + 7s) = 4r + 4t \] Up to here does this make sense?

OpenStudy (bookworm14):

yes that makes sense i was on the right track about multiplying by 6 becuase that gets rid of the fractions all together,

OpenStudy (anonymous):

correct but you got that 2 left on the left. We can do this two ways but this way is simpler. Leave the 2 not multiplied through and remove a 2 on the right \[2(-2r + 7s) = 2(2r + 2t)\] The twos that are pulled out cancel and you get \[-2r + 7s = 2r + 2t\] does this make sense what I did?

OpenStudy (bookworm14):

so in a sense, are you saying if we were to "multiply" the left , then we should "multiply" the right and then even though we do not work it out, we can just cancel them out (hope that made some sense lol)

OpenStudy (anonymous):

yes since they both got 2 in common when i pulled that one out on the right they can cancel. Now to finish the problem... \[-4r = 2t - 7s\] Now we are dividing by a negative again so what I am going to do is divide by just positive 4 and multiply a -1 through the top. \[r = \frac{ -1(2t - 7s) }{ 4 }\] \[r = \frac{ -2t + 7s }{ 4 }\](Choice c)

OpenStudy (bookworm14):

oh i see how you did that ! well i passes that topic fwew i only have 2 topics left should equal out to about 10 probs is there any way you could help me, i promise im not cheating your answers or anything im taking notes and everything !!!

OpenStudy (anonymous):

I can keep helping lol and I believe you!

OpenStudy (bookworm14):

the next one is formulas, im decent at them ,but the more coplex ones i struggle ... fwew Lets do this !!!

OpenStudy (anonymous):

ok so the volume of a sphere is \[V = \frac{ 4 }{ 3 } \pi r^3\] Ad we want to get r by itself. Let me ask you first. What would you do first?

OpenStudy (bookworm14):

Uhm I think i would try to somehow divide "V" by the 4/3 , but im not sure thats right??? (this one is tricky for me)

OpenStudy (anonymous):

well we can divide V by 4/3 yes but an easier way so you dont have two fraction bars is multiply by 3 first then divide by 4 and the pi since pi is just a number. \[V = \frac{ 4 }{ 3 } \pi r^3\] \[3V = 4 \pi r^3\] \[\frac{ 3V }{ 4\pi } = r^3\] Now we need to take a cube root but your answer needs to be in a power form for the root. The rule for roots and powers is  So to take a cube root in power form we essentially take the right and left to the power of 1/3 \[(\frac{ 3V }{ 4\pi })^{1/3} = r\]